Project

General

Profile

NDOS Commissioning Meeting 2011/09/19

Attendance: Jon, Martin, Kurt, Susan, Nick, Rafael, Alex, Enrique, Mat, Denis, Minerba, Gary, Peter

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • NuMI target is to be installed today. Beam should start up again on the 22nd.
  • Shifter last night managed to kill all processes related to startRunControl.sh by killing the terminal where these processes were started. Moving towards iconified, x-window-less start-ups of the DAQ software would help reduce the probability of something like this happening again.

DAQ Status & Plans

  • Kurt: possibly some base release updates (CfgMgr GUI, etc.)
  • Jon: possibly RC and RsrcMgr updates on detector later this week
  • Peter: FTS? FTS has not transferred files since Sept. 13. We'll need to track down Andrew N. to deal with this.

DCS Status & Plans

  • Martin:
    • overview screen on monitor updated to show only FEB temperature alarm status
    • continue to work on getting GUI to control cooling, etc.
    • one more icon will be made available to kill the DCS monitor. Use this if monitor freezes or if one tries to reopen CSS and can't seem to get a session going. Effort in progress to understand why this is happening.

AOB

  • Gary: what's the schedule for APD delivery. Mat: one silicone-coated APD should arrive sometime this week which will be installed and operated cold. Other APDs probably won't be arriving until early Oct.